### v1

# Uncovering Latent Themes of Messaging on Social Media by Integrating LLMs: A Case Study on Climate Campaigns

Abstract
This paper introduces a novel approach to uncovering and analyzing themes in social media messaging. Recognizing the limitations of traditional topic-level analysis, which tends to capture only the overarching patterns, this study emphasizes the need for a finer-grained, theme-focused exploration. Conventional methods of theme discovery, involving manual processes and a human-in-the-loop approach, are valuable but face challenges in scalability, consistency, and resource intensity in terms of time and cost. To address these challenges, we propose a machine-in-the-loop approach that leverages the advanced capabilities of Large Language Models (LLMs). This approach allows for a deeper investigation into the thematic aspects of social media discourse, enabling us to uncover a diverse array of themes, each with unique characteristics and relevance, thereby offering a comprehensive understanding of the nuances present within broader topics. Furthermore, this method efficiently maps the text and the newly discovered themes, enhancing our understanding of the thematic nuances in social media messaging. We employ climate campaigns as a case study and demonstrate that our methodology yields more accurate and interpretable results compared to traditional topic models. Our results not only demonstrate the effectiveness of our approach in uncovering latent themes but also illuminate how these themes are tailored for demographic targeting in social media contexts. Additionally, our work sheds light on the dynamic nature of social media, revealing the shifts in the thematic focus of messaging in response to real-world events.
Introduction
The social media landscape is highly diverse and dynamic, serving as a battleground for various interest groups, including politicians, advertisers, and other stakeholders. These entities strategically use these platforms for microtargeting (Hersh 2015; Barbu 2014), a process that utilizes data-driven techniques to target potential users and advance their interests. Microtargeting capitalizes on the rich user data collected by social networks, making the understanding of this targeted messaging a significant challenge.
In recent years, social media has emerged as a fertile ground for diverse discussions and expressions, encompassing a broad spectrum of topics and themes. In this context, a ‘topic’ within social media content refers to a broad category under which various finer-grained ‘themes’ can be identified. The topic serves as an umbrella term that encompasses a range of more specific and nuanced themes. For example, under the ‘Climate Change’ topic, a theme might be ‘Renewable Energy Solutions’. Identifying themes within a topic involves delving into these specific aspects, revealing the detailed and granular nuances within the broader category. Given the dynamic nature of social media, accurately discerning and understanding the diverse themes within a broader topic is a complex and continuously evolving task.
Traditionally, thematic analysis (TA) has been the method of choice for this task, serving as a qualitative research technique focused on pattern identification, where themes emerging from the data guide further analysis (Roberts, Dowell, and Nie 2019; Vaismoradi et al. 2016; Vaismoradi, Turunen, and Bondas 2013; Braun and Clarke 2012, 2006; Tuckett 2005). However, with the explosion of textual data in the digital age, there’s a growing inclination towards computational methods for text analysis (Sivarajah et al. 2017; Jagadish et al. 2014; Fan, Han, and Liu 2014).
Among these, supervised text analysis methods like Support Vector Machines (SVMs) (Joachims 1998), Naive Bayes classifiers (McCallum, Nigam et al. 1998), and Neural Networks (LeCun, Bengio, and Hinton 2015) have been effective in theme categorization. These methods use pre-labeled datasets to learn and identify patterns, aiding in the categorization of new data. However, they present challenges, notably the need for extensive labeled datasets, which can be resource-intensive to create. Moreover, their performance heavily relies on the quality and representativeness of the training data. If the training data is biased or lacks diversity, the model’s ability to generalize and accurately identify themes in varied or new contexts is significantly compromised.
On the other hand, unsupervised text analysis methods, such as Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003) and non-negative matrix factorization (NMF) (Lee and Seung 1999), have gained popularity in theme discovery. These methods independently identify themes from data without the need for pre-labeled datasets. While useful, they also have limitations. Topic modeling techniques like LDA can produce topics that are challenging to interpret or lack coherence, complicating the extraction of meaningful themes (Röder, Both, and Hinneburg 2015; Lau, Newman, and Baldwin 2014; Mimno et al. 2011; Chang et al. 2009). Recent advancements have seen the integration of human judgment through a human-in-the-loop approach, which, while enhancing accuracy, suffers from issues of scalability and high resource demand (Pacheco et al. 2022a, b; Lund et al. 2017; Hu et al. 2014). Recognizing this, our work explores a new approach, integrating the strengths of LLMs (Brown et al. 2020) with a streamlined algorithmic process to efficiently discover and characterize themes in social media messaging. This exploration leads us to the following research questions (RQ), which are crucial for assessing the potential of LLMs to enhance our analysis by providing deeper and more nuanced insights:
  1. 1.
    RQ1: Can LLMs determine if two given texts, without prior knowledge of existing codes, are discussing the same topic?
  2. 2.
    RQ2: If provided with a definition of a theme, can LLMs successfully categorize other texts under that specific theme?
To enhance theme discovery, we propose an iterative three-step machine-in-the-loop framework. We define large textual collections as repositories of textual instances (e.g., ads, tweets, posts, documents) where each instance is potentially associated with a set of annotated themes (seed set from Islam, Zhang, and Goldwasser (2023) that is used in this paper showed in Table 1). A high-level overview of our method is shown in Figure 1.
Figure 1: Framework overview.
Firstly, we initiate with candidate generation, a phase that includes clustering of instances using a clustering algorithm, ensuring consistency and coherence of the clusters using LLMs, and summarizing each cluster using LLMs. This phase also incorporates LLMs for redundancy checks and merging overlapping clusters, thus refining the cluster pool.
Secondly, we acknowledge the potential need for human validation. This optional phase allows for a critical assessment of the newly formed clusters, questioning their merging, description, and overall coherence. This step serves as a bridge, combining the computational efficiency of algorithms with the nuanced understanding of human judgment.
Finally, the assignment phase utilizes the power of LLMs in a few-shot prompting manner. This innovative approach determines if a new post/text aligns with the summarized clusters, thus assigning it to a relevant cluster.
To showcase our approach, we use climate campaigns (dataset released by Islam, Zhang, and Goldwasser (2023)) on social media as a case study. The dataset focuses on climate related English ads based on the United States from January $2021$ January $2022$ (Islam, Zhang, and Goldwasser 2023). Qualitative researchers use different theories to ground theme discovery, each associated with a different set of concepts. In this case, the theme discovery process is grounded using energy industry and climate change related stances (Islam, Zhang, and Goldwasser 2023). Our contributions are summarized as follows:
  1. 1.
    We suggest a LLMs-in-the-loop approach to uncover latent themes of social media messaging.
  2. 2.
    We conduct quantitative and qualitative analysis on real-world dataset to demonstrate the effectiveness of our proposed method. We show that our model outperforms the baselines.
  3. 3.
    Based on the newly discovered themes, we analyze demographic targeting on social media messaging.
  4. 4.
    We demonstrate how the thematic emphasis of advertisements evolves in response to real-world events.
Pro-energy
Economy_pro, ClimateSolution, Pragmatism, Patriotism, AgainstClimatePolicy.
Clean-energy
Economy_clean, FutureGeneration, Environmental, HumanHealth, Animals, SupportClimatePolicy, AltEnergy, PoliticalAffiliation.
Table 1: Pre-existing themes from Islam, Zhang, and Goldwasser (2023).
Related Work
Several works have done for understanding messaging on social media (Islam, Roy, and Goldwasser 2023; Islam, Zhang, and Goldwasser 2023; Islam and Goldwasser 2022; Capozzi et al. 2021; Silva and Benevenuto 2021; Silva et al. 2020; Serrano et al. 2020; Mejova and Kalimeri 2020; Capozzi et al. 2020; Andreou et al. 2019). Most of the approaches to understanding messaging have relied on broader categories. However, in this paper, we focus on gaining a deeper understanding of messaging.
Previous works (Islam, Roy, and Goldwasser 2023; Islam, Zhang, and Goldwasser 2023; Islam and Goldwasser 2022; Capozzi et al. 2021) used a predefined set of labels, themes, and arguments to analyze messaging. These were fixed and established based on existing topics or theoretical frameworks, such as Moral Foundations Theory (MFT) (Haidt and Graham 2007). However, previous works often fall short of capturing the nuances of messaging choices. For instance, within the Economy topic, it may neglect to recognize conflicting arguments, such as Argument 1: “Increasing the minimum wage boosts economic growth,” versus Argument 2: “Increasing the minimum wage harms small businesses and reduces employment opportunities.” Recently, Pacheco et al. (2022a, b) suggested a human-in-the-loop approach to uncover new themes. In contrast to their work, we use LLMs inference.
Researchers in the fields of qualitative research (QR) and Natural Language Processing (NLP) are currently investigating how LLMs can be utilized for TA. Dai, Xiong, and Ku (2023) proposed LLM-in-the-loop approach for TA. Xiao et al. (2023) combined GPT-3 with expert-drafted codebooks for supporting qualitative analysis. Gao et al. (2023) suggested a collaborative interface for qualitative analysis in their work, utilizing LLMs for generating initial codes and aiding in decision-making processes. In this paper, we develop a new set of themes (or codes) specifically tailored for analyzing messaging in a particular domain, which helps speed up the development of domain-specific labels.
Framework
Our methodology is meticulously designed to discover themes in social media messaging through a robust three-step process. This approach not only capitalizes on advanced computational techniques but also integrates human insight, ensuring a comprehensive and accurate analysis. Importantly, our framework is flexible regarding the choice of representation. Both the embedding technique used for generating representations and the scoring function can be readily substituted with alternatives.
Theme and Instance Representation
We represent instances and explanatory phrases under each theme using the Sentence BERT (Reimers and Gurevych 2019) embedding. To evaluate the degree of relevance between an instance and a specific theme, we adopt a straightforward distance-based method. This involves calculating the cosine distance between the instance and every explanatory phrase and example related to the themes, selecting the smallest distance score as the representative value. Our framework is agnostic of the representation used. The underlying embedding objective, as well as the scoring function, can easily be replaced.
Grounding
Before assigning instances to themes, every instance is grouped into unnamed clusters, effectively leaving them unassigned. Upon the first invocation of this process, instances are assigned to the theme that is nearest to them, but only if the newly calculated distance is shorter than, or at most equal to, the distance from their prior assignment. Previous assignments can correspond either to different themes or to the unnamed space. Note that this way, some instances can remain unassigned based on a threshold.
Figure 2: Example of incoherent cluster.
Candidate Generation
The initial phase of our methodology, candidate generation, begins with the clustering of unassigned instances. We employ K-means clustering (Jin and Han 2010), a method chosen for its effectiveness in handling large datasets, which consist of varied and complex social media messaging. The K-means algorithm segments these instances into distinct clusters based on their thematic and linguistic similarities.
Figure 3: Example of merged cluster.
Coherency check
To ensure the consistency and coherence of these clusters, we perform a cluster coherence check. We create subsets of instances from each cluster, with each subset containing $4$-$5$ instances. We then prompt LLMs in a zero-shot setting, asking if these instances belong together. Based on the responses provided by the LLMs, we identify incoherent clusters as noise and exclude them. It is important to note that we discard only the clusters, not the instances associated with them. These instances are still utilized for cluster membership assignments in the subsequent stage of our framework. Fig. 2 shows an example of an incoherent cluster from the dataset. These are the top $5$ ads that are closest to the centroid of a cluster. $1^{st}$ and $2^{nd}$ ads are about supporting motor oil, $3^{rd}$ one is against the oil and gas industry, $4^{th}$ and $5^{th}$ ads are about supporting biofuel. LLMs categorize this cluster as an incoherent cluster (answering RQ1).
Cluster Summarization
To make the clusters interpretable, we prompt LLMs in a zero-shot setting to generate a multi-document summary based on top-$K$ instances of the clusters ($k=5$).
Redundancy check
We embed the summaries using SBERT (Reimers and Gurevych 2019) and compute the cosine similarity score between the embedding summaries of cluster pairs. Based on a threshold value ($\geq 0.6$), we merge two clusters. Fig. 3 shows an example of a merged cluster from the data. These cluster summaries are related to “initiatives to plant new trees and protect existing forests to save planet” and based on the threshold, we merge them.
Figure 4: Prompt template for coherency check (shown as zero-shot)
Figure 5: Prompt template for generating summary (shown as zero-shot))
Figure 6: Prompt template for LLMs Mapper (shown as four-shot)
Human Validation
The second phase introduces an optional but crucial step – human validation. Two NLP and Computational Social Science (CSS) researchers did one $2$ hours session to assess the clusters formed by the algorithms. Their expertise is particularly invaluable in evaluating the nuances of language and context that automated systems might overlook. They scrutinize the merging decisions, the descriptiveness of the clusters, and overall coherence, bridging the gap between computational efficiency and the nuanced understanding of human judgment.
For the cluster coherency check, the annotators are asked to verify if the top $5$ ads from a cluster are coherent or not. They are asked to provide a score of $1$ if the ads are coherent and a score of 00 if they are not.
For the cluster summary check, the annotators are asked to verify if the generated summary is correct or not. They are asked to provide a score of $1$ if the summaries are correct and a score of 00 if they are not.
Map Instances to Themes
In the final phase, we provide summaries of all clusters (including pre-existing + newly discovered) and prompt LLMs in a four-shot setting to assess whether a new instance aligns with the summarized clusters. To alleviate the issues of being hallucinated, we force LLMs to answer the questions by selecting from the given summarized clusters. The LLMs, equipped with the context from the initial phases, effectively categorize new texts into the relevant clusters (answering RQ2). This phase is crucial for maintaining the dynamism of the clusters, allowing the system to adapt to new data and emerging themes.
Method Num. Num. Covered Ads
Themes thr $<0.6$ thr $<0.5$ thr $<0.4$ thr $<0.3$
Pre-existing 13 14652 9725 3731 558
+After Iter1 20 18702 14583 8646 2944
+After Iter2 25 18988 15052 9079 3180
Table 2: Coverage. THR: Threshold, NUM: Number.
(a) Assignment with LLMs mapper.
(b) Assignment with SBERT.
Figure 7: Confusion Matrix after $2^{nd}$ round of iteration + assignment.
(a) Baseline: 10 LDA Topics.
(b) Baseline: 30 LDA Topics.
(c) Ours: After $2^{nd}$ round of iteration.
Figure 8: Correlations between themes and stances.
Experiments
We use $21372$ ads and $13$ themes (Table 1) related to climate campaigns released by Islam, Zhang, and Goldwasser (2023). At first, we assign pre-existing themes to the ads using SBERT. Pre-existing themes can cover a small portion of ads ($1^{st}$ row of Table 2). The threshold is determined by the distance from the centroid to an ad, with a shorter distance being preferable. For a threshold of $<0.5$, there are $11647$ ads that remained uncovered, which we consider as a noisy set. We then apply k-means clustering (with $k=29$) to this noisy set. Following the candidate generation phase, we identify $7$ new themes in the $1^{st}$ iteration. We use GPT-3.5 for cluster coherency check, generating cluster summaries, and redundancy check. In $2^{nd}$ iteration, we repeat this process using $6789$ ads remaining from iteration 1 (again applying a threshold of $<0.5$), resulting in the discovery of $5$ additional themes. The complete set of final themes is presented in Table 3. Summaries of $25$ themes (clusters) are provided in Appendix A.
Iter1
BidenGasPriceIncrease, AgainstCorporateInterests, GasTax, Deforestation, Carbon, CustomerBasedAltEnergy, FoodSecurity
Iter2
EnergyAffordabilityandSustainabilityLegislation, EcofriendlyConsumerChoices, PlasticWasteandEnvironmentalImpact, PromoteSustainableTransportation, WaterManagementandSustainability
Table 3: Resulting Themes. Iter: Iteration.
Method Acc. Macro avg. F1
SBERT Assignment 84.05 79.32
LLMs Mapper (Ours) 88.15 89.24
Table 4: Mapping Quality w.r.t Human Judgements
Prompt Templates
This section shows the prompt templates used to check cluster coherency (Fig. 4), generate cluster summary (Fig. 5), and map ads to themes (Fig. 6).
Age Group
State
Entity
Claim
Theme
13-24 CA
Young People, Climate Change Denier, Climate Movement.
Importance of action against climate change to clean air & water and to reduce carbon emissions.
Environmental
TX
Natural Gas Systems, Coal Energy Systems, Nuclear Energy Systems, Wind Turbines, Solar Panels.
Failures in natural gas, coal, and nuclear systems caused almost double outages in Texas compared to frozen wind turbines and solar panels.
AltEnergy
25-54 CA
Carbon Emissions, Sacramento Municipal Utility District.
Eliminating all carbon emissions from the power supply by 2030.
Carbon
TX
Climate Change, Renewable Energy Plan.
Climate change impacts are irreversible by 2030, urging an immediate switch to renewable energy.
AltEnergy
55+ CA
Western Burrowing Owls, Aramis Industrial Solar Power Plant.
Aramis solar plant’s construction will harm wildlife and habitats, especially affecting Western burrowing owls.
Animals
TX
U.S. Senators or Congressman, Citizens’ Climate Lobby.
Urgent need for action on climate change.
SupClimPolicy
Table 5: Most mentioned entities, claim, and theme of targeted ads for young, working-age, older population of CA and TX. SupClimPolicy: SupportClimatePolicy.
Results and Analysis
Our experiments show that our framework can be used to uncover a set of themes that cover a large portion of the messaging regarding climate campaigns (Table 2). With a threshold $<0.4$, pre-existing themes can cover only $17.5\%$ ads. After $1^{st}$ iteration using our machine-in-loop approach, we cover $40.5\%$ and after $2^{nd}$ iteration, we cover $42.5\%$ ads (Table 2).
Gender
Entity
Claim
Theme
Male
Alex Askew, Karen Greenhalgh
Alex Askew’s record of raising taxes contrasted with Karen Greenhalgh’s promise to cut taxes and lower living costs.
GasTax
Female
Parents, Children/Kids
Urgent need for action against climate change to protect the future of children and the planet.
FutureGen
Table 6: Mentioned entities, claim, and theme of targeted ads for only male and only female population. FutureGen: FutureGeneration.
State Stance
Entity
Claim
Theme
VT Clean-energy Community
Climate change impacts health, family, and community.
HumanHealth
Pro-energy
Climate Action Council
Climate Action Council plans to increase fuel and carbon taxes.
GasTax
ND Pro-energy Community Oil and gas supports jobs and community growth. Economy_pro
Biden
Biden’s energy policies increase energy costs.
BidenGasPrInc
Clean-energy
Water system
Solution of water scarcity problem.
WaterMng&Sustain
Table 7: Entity, claim, and theme of targeted ads for blue and red states. The entity ‘Community’ is targeted differently by advertisers. While targeting the blue state Vermont, the emphasis is on the impacts of climate change on community health. In contrast, when targeting the red state North Dakota, the focus shifts to how oil and gas supports community growth. BidenGasPrInc: BidenGasPriceIncrease, WaterMng&Sustain: WaterManagementandSustainability.
Mapping Quality
To evaluate the mapping quality of our model, we select ads that fall under threshold $<0.4$ and then randomly sample $1072$ ads. We ask the annotators to provide a score of $1$ if the ad aligns with the theme, 00 for no alignment. There is no clear definition of alignment, and we let the annotators make this judgment. To evaluate mapping quality, we compare the performance of LLMs mapper with SBERT assignment (Baseline) of ads to themes. Results are outlined in Table 4. The resulting mapping from ads to the theme is fairly accurate with respect to human judgments (Fig. 6(a)) with the accuracy of $88.15\%$ and macro avg. F1 score of $89.24\%$. Fig. 7 shows the confusion matrix after $2^{nd}$ round of iteration and assignment with LLMs mapper (6(a)) and SBERT mapper (6(b)).
Qualitative Analysis
Given that we can characterize themes as the reasons cited by people to support climate debate, we consider assignments to be better if they are more cohesive (e.g., if they are more strongly correlated with climate-related stance). To evaluate this, we perform a correlation test between the identified themes and the stance expressed in the ads (i.e., pro-energy or clean-energy). We calculate the Pearson correlation matrices and present them in Fig. 8. We compare our discovered themes with a set of topics extracted using LDA. We observe that our themes (Fig. 7(c)) have stronger correlations with stances than the derived LDA topics (Fig. 7(a),7(b)).
Hyperparameters
To obtain topics from LDA, we use Gensim (Rehurek and Sojka 2011) implementation. We follow the preprocessing steps shown in Hoyle et al. (2021) and estimate the number of topics in a data-driven manner by maximizing. We do a grid search over a set of {10, 20, 30} for the LDA baseline.
To determine the value of $k$ in k-Means, we follow both The Elbow Method and The Silhouette Method. Our optimal $k$ value is $29$.
For SBERT embedding, we use the sentence transformer model ‘all-mpnet-base-v2’ with default parameters.
(a) Exxon Mobil.
(b) American Petroleum Institute.
(c) Consumer Energy Alliance.
(d) Apex clean energy.
(e) American Wind Energy Association.
(f) Climate Power Education Fund.
Figure 9: Pro-energy and clean-energy Themes differ based on regions. Funding entities are categorized based on Corporation, Industry Association, and Advocacy Group.
(a) Event: Gas price increase on 12 May, 2021.
(b) Event: Hurricane Ida hits on 29 August, 2021.
Figure 10: Advertisers’ use of the top-$5$ themes before and after the events. (a) Themes ‘GasTax’ and ‘BidenGasPriceIncrease’ emerged mainly after the occurrence of the event Gas Price Increase. (b) Theme ‘SupportClimatePolicy’ emerged after the event Hurricane Ida hits.
Demographic Targeting
Analyzing demographic targeting in social media messaging is crucial for understanding the varying impacts of targeted content on different population segments. This analysis is vital for comprehensively understanding how messaging is tailored and disseminated to influence specific demographic groups.
How Themes Differ based on Age Group
In the dataset (released by Islam, Zhang, and Goldwasser (2023)), there are $7$ age groups. In this paper, we categorize these into three distinct age groups: young people (ages $13$-$24$), working-age people (ages $25$-$54$), and the older population (age $55+$). We analyze the targeted ads for these $3$ age categories, focusing on the states of California (CA) and Texas (TX). The reason for choosing these two states is that TX is known as the energy capital of the world***https://www.eia.gov/todayinenergy/detail.php?id=49356. Recently, CA has emerged as one of the most vocal states in the fight against climate changehttps://www.pewtrusts.org/en/research-and-analysis/blogs/stateline/2022/10/06/california-takes-leading-edge-on-climate-laws-others-could-follow. To identify the top-k most frequently mentioned entities in the texts, we prompt LLMs in a zero-shot setting. Similarly to extract the most mentioned claim, zero-shot prompting of LLMs has been used.
Table 5 summarizes the results. Ads targeted for young population of CA focus on ‘Environmental’ theme having claim ‘Importance of action against climate change to clean air and water’. For the working-age demographic in CA, the ads focus on the ‘Carbon’ theme, making claims like ‘The Sacramento Municipal Utility District (SMUD) aims to eliminate all carbon emissions from their power supply by 2030, contributing to the #CleanPowerCity movement’. In TX, both young and working-age groups are the focus of ads under the ‘AltEnergy’ theme, each with specific claims. Interestingly, the older population in CA is targeted by ads with claims opposing the construction of solar plants, such as The construction of solar plants will harm wildlife and habitats’, aligning with the ‘Animals’ theme. This approach is interesting, as it positions advertisers against solar plant construction while still engaging with the broader context of clean-energy theme.
How Themes Differ based on Gender
To analyze how messaging is tailored for a specific gender, we extract ads targeted exclusively at males and females, focusing on the themes that are uncommon between these categories. We find that ads having ‘GasTax’ theme focus only male population. On the other hand, ‘FutureGeneration’ themed ads only target female population with mentioned entities ‘Parents’, ‘Children/Kids’ and the claim emphasizing the ‘Urgency of taking action against climate change to protect children’s future’ (Table 6).
How Themes Differ State-wise
To explore the state-wise differences in emerging themes and claims, we use historical voting datahttps://www.visualcapitalist.com/u-s-presidential-voting-history-by-state/ to categorize states as red or blue. We focus on Vermont (VT), a blue state, and North Dakota (ND), a red state. A notable finding is the different emphasis placed on the entity ‘Community’ by VT and ND. In VT, the focus is on the claim Climate change impacts health, family, and community’ (under the theme ‘HumanHealth’), while in ND, the emphasis is on Oil and gas support jobs and community growth’ (falling under the theme ‘Economy_pro’). These contrasting perspectives are detailed in Table 7. Besides, ads having pro-energy stance focus on ‘GasTax’ theme while targeting blue state. Conversely, ads having clean-energy stance focus on ‘WaterManagementandSustainability’ theme while targeting red state (Table 7).
Pro-energy
Corporation: Exxon Mobil, Industry Association: American Petroleum Institute, Advocacy Group: Consumer Energy Alliance.
Clean-energy
Corporation: Apex clean energy, Industry Association: American Wind Energy Association, Advocacy Group: Climate Power Education Fund.
Table 8: List of funding entities.
How Themes Differ based on Geographic Regions
To grasp the ways in which fossil fuel industries and their supporting groups shape public opinion, Islam, Zhang, and Goldwasser (2023) categorize pro-energy funding entities into three distinct types: Corporation, Industry Association, and Advocacy Group. Similarly, we categorize the clean-energy funding entities into those categories. In this paper, we choose only one funding entity from each category resulting $6$ funding entities, including $3$ pro-energy and $3$ clean-energy funding entities. Table 8 shows the list of funding entities included in our analysis.
We divide the USA into $5$ regions, encompassing all $50$ states: West, Southwest, Midwest, Southeast, and Northeast to analyze the geographic targeting of the sponsored ads (Fig. 9). Exxon mobile targets Northeast region sponsoring ads focused on ‘energy affordability and reliability’ (‘Pragmatism’ theme). For the rest of the regions, it sponsor ads focusing on ‘how energy can play in supporting jobs, wages and community growth’ (‘Economy_pro’ theme). On the other hand, the pro-energy advocacy group Consumer Energy Alliance focuses on the ‘Economy_pro’ theme when targeting the Midwest, but shifts to the ‘ClimateSolution’ theme for the Southeast region (Fig. 8(c)).
From Fig. 8(d), we notice that Apex clean energy targets the West region focusing on ‘EcofriendlyConsumerChoices’ and ‘PromoteSustainableTransportation’ themes, while for other regions (except the Southwest) it emphasizes on ‘CustomerBasedAltEnergy’ theme. However, American Wind Energy Association uses ‘PoliticalAffiliation’ theme to target the Southwest and Midwest regions (Fig. 8(e)).
Theme Shifts Triggered by Key Events
To investigate how theme differs based on real-world event, we pick 2 defining events and investigate the changes of the themes. The events are as follow:
The usage of themes in response to these events are shown in Fig. 10. Fig. 9(a) illustrates a comparative bar chart of the sponsored ad themes (top-$5$) used five days before and after a significant event, labeled as the Gas Price Increase. We observe that the frequency of themes ‘AltEnergy’, ‘AgainstCorporateInterests’, and ‘ClimateSolution’ increased following the event. Additionally, the themes ‘GasTax’ and ‘BidenGasPriceIncrease’ emerged predominantly in the aftermath (Fig. 9(a)).
Fig. 9(b) shows the shift in the thematic focus of sponsored advertisements over a period of six months before and after Hurricane Ida hits in the USA. The emerging themes after the hurricane, such as ‘SupportClimatePolicy’, reflect a growing public and corporate response to environmental issues highlighted by climate change. There is a noticeable presence of themes related to alternative energy (Fig. 9(b)).
Overall, Fig. 10 shows a shift in the thematic focus of ads in response to the real-world event.
Conclusion
We present a machine-in-the-loop framework for discovering new themes from social media messaging in a particular domain, which can help speed up the development of domain-specific labels. We show that themes generated from our approach can cover a larger portion of ads, and those themes are coherent through quantitative measures and human evaluation. Additionally, we demonstrate that our resulting set of themes is cleaner and more explainable than themes obtained with topic modeling approaches.
Furthermore, our analysis demonstrates how these themes are specifically tailored for demographic targeting on social media and how the thematic focus of messaging shifts in response to real-world events.
Limitations
Though we show the effectiveness of our framework focusing on climate campaigns, it is a domain-agnostic framework. However, the main idea of characterizing themes using LLMs and associating the ads to their closest themes is applicable to other cases with no changes.
Our approach relies on GPT-3.5 to generate a multi-document summary, check cluster coherency, and map ads to themes. We chose GPT-3.5 instead of the open-source counterparts due to computational resource constraints.
LLMs are pre-trained on a huge amount of human-generated text. As a result, they may inherently contain many human biases (Blodgett et al. 2020; Brown et al. 2020). We did not consider any bias in our task.
Ethics Statement
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details needed for reproducing the results.
In this paper, we did not propose any new dataset rather, we only experiment with an existing dataset which are adequately cited.
The author’s personal views are not represented in any qualitative result we report, as it is solely an outcome derived from a machine learning model.
Broader Impact
Our approach marks a significant advancement in the thematic analysis of social media messaging, which might provide valuable insights into the strategic use of messaging for targeted communication and campaign efficacy. We leave this exploration for future work.
Appendix A Appendix
Cluster Summaries
Summaries of final $25$ themes (clusters) are shown below:
1. Summary_Economy_pro: The provided texts collectively emphasize the importance of the oil and natural gas industry in supporting local jobs, economic growth, and community development. They express concerns about proposed bans on natural gas and oil, warning of potential job losses and negative economic impacts. The texts call for support of the oil and gas industry, highlighting its contributions to local economies and communities while advocating for a sustainable future.
2. Summary_ClimateSolution: The provided texts collectively advocate for the role of natural gas in advancing a cleaner and more sustainable energy future. They highlight natural gas as an essential tool for reducing carbon emissions, providing affordable and reliable energy, and complementing renewable energy sources like wind and solar. The texts emphasize the importance of using natural gas to address the dual challenge of meeting energy needs and protecting the environment while working towards ambitious carbon reduction goals, such as reaching net-zero emissions by 2045. Texts focus on greenwashing aspect and the promotion of a controversial energy source under the guise of environmental benefit. The text covers the complex and sometimes contradictory narratives around certain fossil fuels in the context of climate and energy policy.
3. Summary_Pragmatism: The provided texts collectively stress the vital role of reliable and affordable natural gas and oil in enabling modern life. They highlight that in 2020, natural gas and oil accounted for nearly $70\%$ of America’s energy supply, and projections suggest they will continue to meet approximately $70\%$ of the nation’s energy needs by 2050, as indicated by the U.S. Energy Information Administration. The texts call attention to the significance of American oil and gas production in maintaining affordable energy and invite individuals to share their priorities in this regard.
4. Summary_Patriotism: The provided texts collectively emphasize the importance of U.S. energy leadership and innovation, highlighting the country’s position as a top global energy producer while actively working to reduce greenhouse gas emissions. They express concerns about potential cutbacks in local oil and gas production, citing potential consequences such as increased reliance on foreign oil imports and higher gas prices for Californians. The texts argue against shutting off American oil and natural gas production, pointing out that such actions could strengthen other nations while negatively impacting U.S. jobs, economic security, and national security. Additionally, they underscore the broader impacts of U.S. oil and natural gas production on daily life, including its role in fertilizer production and its significance in ensuring food supply and availability. The texts call for action to protect America’s energy future and support natural gas and consumer choice through petitions, and they highlight the U.S.’s achievement as the world’s No. 1 exporter of liquefied natural gas, particularly in response to energy demands in Europe.
5. Summary_AgainstClimatePolicy: The provided texts collectively convey a sense of urgency and concern regarding the oil and gas industry, emphasizing the need for support and resistance against what they perceive as threats to the industry’s survival. They criticize progressive and Democratic energy policies, particularly the Green New Deal, as potentially devastating to American workers and the economy. The texts also criticize the Biden administration’s energy policies, linking them to high gas prices and a reliance on emergency reserves. They call for action to prioritize and protect the American energy industry, with an emphasis on opposing what they view as radical policies and mandates. Additionally, the texts highlight the role of the oil and gas industry in addressing climate challenges through innovation and expertise while expressing fears of job loss and economic consequences if certain policies are implemented.
6. Summary_Economy_clean: The provided texts collectively advocate for a clean energy future in the United States and emphasize the potential for job creation and economic benefits associated with clean energy initiatives. They highlight the growth of clean energy sectors and argue that investments in renewable energy are generating good-paying jobs across various occupations and regions. The texts contrast clean energy with fossil fuels, suggesting that clean energy is not only environmentally responsible but also economically advantageous. They call on Congress to support clean energy policies and investments as a means to strengthen the American economy, create jobs, lower energy costs, and mitigate climate change, ultimately promoting a clean energy-focused approach to economic growth.
7. Summary_HumanHealth: The provided texts collectively emphasize the significant and urgent health impacts of climate change, highlighting the range of dangers it poses, including cardiovascular and respiratory issues, food security, the spread of diseases, and more. They stress the need for awareness and action to address these health challenges linked to climate change, with mentions of governmental initiatives and organizations dedicated to addressing climate-related health concerns. The texts underscore that climate change is already harming people and communities, affecting vulnerable populations, and call for support and action to mitigate these adverse health effects.
8. Summary_FutureGeneration: The provided texts collectively emphasize the grave threat of climate change to children’s health and well-being. They cite alarming statistics, including data from the World Health Organization, that underscore the vulnerability of young children to climate-related health issues. The texts stress the urgent need to protect children from the impacts of climate change and highlight the various health risks they face, including premature births, birth defects, and increased emergency room visits. They call for action to address climate change to safeguard the future of children and grandchildren, emphasizing that inaction comes at a high cost to children’s health and well-being. The texts also acknowledge the anxiety and concerns that climate change is causing among young people and suggest ways to support and address their anxieties.
9. Summary_Environmental: The provided texts collectively address the urgency and breadth of the climate crisis, emphasizing its far-reaching impacts on various aspects of society and the environment. They highlight alarming trends such as the rapid warming of oceans, the increasing frequency of extreme weather events, and the dire consequences of climate change on ecosystems, economies, and public health. The texts call for collective action and emphasize the role of individuals in combating climate change, urging the sharing of ideas and solutions. They also stress the need to address the root causes, including greenhouse gas emissions from fossil fuels, and promote carbon capture as a key strategy for decarbonizing the global energy system. Overall, the texts underscore the severity of the climate emergency and the importance of taking immediate action to mitigate its effects and protect vulnerable communities and ecosystems.
10. Summary_Animals: The provided texts collectively highlight the severe impact of climate change on wildlife, fisheries, birds and ecosystems. They mention rising temperatures, increased droughts, fires, intense storms, and floods that threaten various species and their habitats. Studies reveal that climate change could lead to the extinction of one-third of species in the next 50 years, with particular vulnerabilities in South America. The texts emphasize the urgent need for action to protect wildlife from the consequences of climate change, emphasizing the role of renewable energy development that must consider the needs of wildlife. They call for an understanding of the interconnected crises of climate change and biodiversity loss, urging collective efforts to address both challenges.
11. Summary_AltEnergy: The provided texts collectively advocate for the adoption of renewable energy sources as a cleaner, cheaper, and more sustainable alternative to fossil fuels. They emphasize the benefits of transitioning to renewables, including reducing carbon footprints, enhancing grid reliability, and mitigating the impact of rising gas prices. The texts highlight the rapid growth of renewable energy in the United States, with clean power sources like wind, solar, and hydrogen accounting for a significant share of electricity generation. They stress the urgency of making this transition to combat climate change, citing the irreversible impacts looming by 2030. Additionally, they mention various projects and initiatives aimed at promoting clean energy and energy independence, such as repurposing coal plants into renewable energy hubs and proposed solar farms. Finally, they encourage energy-conscious practices like using electricity during off-peak hours to harness clean energy effectively.
12. Summary_SupportClimatePolicy: The provided texts collectively advocate for the passage of the Build Back Better Act, emphasizing its significance in addressing climate change and promoting a clean energy future. They highlight the key provisions of the bill, including cutting pollution by $50\%$ by 2030, creating clean energy jobs, and lowering utility bills. The texts stress the urgency of taking action to combat climate change and emphasize that it’s a historic opportunity to invest in meaningful climate action and transition to a carbon-zero grid. Additionally, some texts criticize the fossil fuel industry for its role in climate change and call for holding polluters accountable. Overall, the focus is on the importance of the Build Back Better Act in advancing climate goals and environmental justice.
13. Summary_PoliticalAffiliation: The provided texts collectively highlight concerns about the influence of the oil and gas industry on climate change policies. They emphasize that despite claims of caring about climate change, these industries spend significant amounts on lobbying to block regulations and protect their subsidies. The texts mention the substantial sums spent on climate lobbying by major oil and gas companies and argue that these actions hinder meaningful climate action. They also discuss political contributions to lawmakers and emphasize the need to address climate change without being influenced by industry lobbying. Additionally, some texts reference specific instances where lobbyists have allegedly undermined climate legislation or influenced infrastructure bills to prioritize industry interests over environmental concerns. Overall, the focus is on revealing the industry’s efforts to shape climate policies to their advantage.
14. Summary_BidenGasPriceIncrease: The provided texts collectively express strong disapproval and criticism of President Joe Biden’s administration, focusing on various issues such as rising gas prices, inflation, the southern border crisis, and energy policies. Some texts call for Joe Biden’s resignation and invite readers to participate in approval polls, while others highlight specific concerns, like the increase in gas prices and the cancellation of certain pipelines, portraying these actions as detrimental to the American economy. Overall, these texts present a negative view of Joe Biden’s leadership and policy decisions.
15. Summary_AgainstCorporateInterests: The provided texts largely revolve around progressive candidates running for Congress and their calls for grassroots support. These candidates emphasize their commitment to economic justice, racial equity, climate action, and progressive policies such as a Green New Deal, Medicare for All, and universal housing. They criticize their opponents for receiving significant contributions from corporations and lobbyists and highlight the need for people-powered campaigns. Some texts also mention the importance of supporting incumbent progressive Congress members, such as the Squad, to pass transformative legislation like the Build Back Better Agenda. Overall, these texts aim to rally support for progressive candidates and causes while advocating for a shift away from corporate influence in politics.
16. Summary_GasTax: The provided texts revolve around opposing proposed gas tax increases in various states. They emphasize concerns about the financial burden on families, rising gas prices, and the potential for gasoline shortages. The texts encourage people to sign petitions or participate in rallies to voice their opposition to these tax hikes. Additionally, some texts highlight the current tax rates and the impact of previous tax increases, framing the issue as a matter of transparency and affordability for consumers. Overall, the focus is on mobilizing public sentiment against gas tax increases in different states.
17. Summary_Deforestation: The texts primarily focus on environmental conservation and efforts to combat climate change. They highlight the importance of preserving forests like the Tongass National Forest in Alaska, emphasizing their role in storing carbon, providing habitat, and mitigating climate change effects. They urge readers to take action, such as signing petitions or participating in tree-planting initiatives. Some texts discuss the potential benefits of ”super trees” that are more effective at carbon storage and stress the importance of preserving mature trees in the face of climate change. Overall, the texts call for collective action to protect and restore natural environments and address climate-related challenges by planting more trees and preserving forests.
18. Summary_Carbon: These texts collectively emphasize the importance of reducing carbon emissions and mitigating one’s carbon footprint. They introduce various strategies and initiatives for individuals and communities to lower their carbon impact, such as carbon pricing, using energy-efficient lights, enrolling in carbon reduction programs, calculating personal carbon footprints, offsetting tailpipe emissions, and supporting sustainability efforts in industries like cement and concrete. The overarching theme is the need for environmental awareness and action to combat climate change and promote a more sustainable future.
19. Summary_CustomerBasedAltEnergy: These texts collectively highlight various aspects of renewable and sustainable energy sources. They address common myths and misconceptions about solar energy, emphasizing its versatility and suitability for different types of buildings and applications. They promote community solar programs in Illinois, offer information and resources for learning about solar power, and invite individuals to explore solar technology and installations. Additionally, the texts touch upon other forms of renewable energy, such as wind and nuclear power, showcasing their potential to provide reliable, clean energy and create jobs in various states. The overarching theme is the promotion and education about renewable energy sources and their benefits.
20. Summary_FoodSecurity: The provided texts collectively address the complexities of food security, emphasizing the global nature of the challenge and the need for cross-sector collaboration to meet the increasing food demand caused by a growing global population and climate change. They also shed light on issues of food justice and sovereignty, particularly in marginalized communities, where access to healthy foods is limited. Additionally, the texts discuss the concept of food elitism, questioning marketing practices and their impact on food equity, arguing for the importance of affordable and sustainable food systems. Furthermore, the issue of food deserts is explored, where residents face challenges related to the availability, accessibility, and affordability of nutritious food, with a recognition of the intricate factors contributing to food insecurity. These texts collectively stress the urgency of addressing food-related challenges from multiple perspectives, encompassing policy, social justice, and sustainable agriculture.
21. Summary_EnergyAffordabilityandSustainabilityLegislation: These texts all convey similar messages, focusing on various representatives, including Rep. John Autry, Rep. Kelly Alexander, Rep. Kyle Hall, Rep. John Sauls, and Rep. John A. Torbett, who are collectively working to reduce power bills for families and businesses in North Carolina. They sponsor legislation aimed at exploring changes that could potentially result in annual energy bill reductions exceeding \$$\$600$ million. The texts encourage readers to learn more about the Alliance for Energy Security’s collaboration with policymakers in an effort to make energy in North Carolina more affordable, environmentally friendly, and dependable.
22. Summary_EcofriendlyConsumerChoices: The broader theme of these texts is environmental sustainability and eco-conscious consumer choices. They focus on promoting a laundry detergent product that aligns with sustainability goals by offering a range of environmentally friendly features, such as being cruelty-free, using minimal plastic packaging, and supporting carbon-neutral shipping. The theme encourages individuals to make environmentally responsible choices in their daily lives, particularly in the context of laundry and reducing plastic waste.
23. Summary_PlasticWasteandEnvironmentalImpact: The overall summary of these texts is a collective call to address the urgent issue of plastic waste and pollution in the environment. They highlight the alarming amount of plastic entering our oceans, stress the need for collective action involving both consumers and brands to combat this crisis, introduce innovative solutions like pyrolysis to reduce plastic waste and carbon footprints, and emphasize the importance of better handling and regulation within the petrochemical industry to prevent plastic pollution. The texts underscore the broader theme of environmental responsibility and the imperative to find sustainable alternatives to single-use plastics and improve the entire plastic production and waste management lifecycle.
24. Summary_PromoteSustainableTransportation: These texts collectively emphasize the importance of electric vehicles (EVs) in combating air pollution and improving air quality. They highlight the role of clean car standards and the accessibility of EVs, particularly in Los Angeles, where local rebate programs make them a cost-effective choice. Shifting from gas-powered vehicles to electric options is promoted for its potential to reduce emissions and decrease exposure to pollution, with a focus on the environmental and health benefits of adopting EVs and the need to take steps to clean the air in Colorado and beyond.
25. Summary_WaterManagementandSustainability: These texts collectively emphasize the discussion of water treatment and billing issues in various cities. The texts touch upon topics such as the cost of building new water treatment plants versus refurbishing existing ones, concerns about water quality and the presence of toxic chemicals in tap water, allegations of potential corruption or kickbacks in decision-making processes, and the financial aspects of water billing, including rates, fees, and expenditures related to water infrastructure. Additionally, some texts emphasize the importance of proper maintenance and infrastructure investment to ensure clean and safe drinking water for residents.